OpenStudy (anonymous):

How would I graph the Parabola on a parent graph? y=(x+3)^2

OpenStudy (anonymous):

move the graph to the left 3

OpenStudy (anonymous):

Left 3 and up 3?

OpenStudy (anonymous):

no just left three

OpenStudy (anonymous):

the equation you have is already in vertex form, y=(x+3)^2 or better said as y=1(x-(-3))^2 + 0 [do you see how it is still the same equation] this is now in a similar form to the standard vertex form of y=a(x-h)^2 + k where (h,k) are the values of the vertex of the parabola